OpenStudy (anonymous):

can you simplify (√3)^3

OpenStudy (anonymous):

yes i can...

OpenStudy (anonymous):

can you help me

OpenStudy (karatechopper):

well first lets expand it!

OpenStudy (karatechopper):

\[\sqrt{3}\times \sqrt{3}\times \sqrt{3}\]

OpenStudy (karatechopper):

now can u solve? or need more help

OpenStudy (anonymous):

all you need to do now is simplify what karatechopper did..

OpenStudy (anonymous):

@karatechopper is the answer 3√3

OpenStudy (karatechopper):

@dpaInc thanks:) now dont say anything now im quizzing... @iHeartYou idk do you think it is?

OpenStudy (anonymous):

@karatechopper yeah

OpenStudy (karatechopper):

then thats your answer

OpenStudy (anonymous):

thanks

OpenStudy (karatechopper):

welcome!